What is the y-intercept of the graph of the equation
6x − 5y + 15 = 0?

Answer by ikleyn(52788)

.

y-intercept is the value of  "y"  from the equation, when x= 0.


In your case, y-intercept is


    -5y + 15 = 0

    15  = 5y

     y  = 15/5 = 3.    ANSWER

Solved, answered and explained.

                M E M O R I Z E !


            y-intercept is the value of  " y "  from the equation,  when x= 0.